# Measuring the (Un)Faithfulness of Concept-Based Explanations

### v2

@@ 4.1 Deletion-Based Measures
Deletion Strategy: defines how concepts are removed. Concepts are usually removed from the Perturbation Space by setting the concept to a baseline value (e.g., 0).
Y-axis: defines what is measured after each deletion. Typically, this is accuracy (e.g., in object classification).
X-axis: represents how much “information” was deleted. This is typically the number of deleted concepts.
@@
To understand the differences between prior deletion measures, Tab. 1 organizes prior measures according to these dimensions. ACE [13] deletes concepts in the image space by setting associated pixels to a constant value (Constant Deletion). They measure the model’s prediction accuracy after each deletion across a dataset. CRAFT [9] transforms the model’s embeddings into a concept representation. It makes Constant Deletions in the concept space by setting the concept to zero; then, they reconstruct the embedding from the perturbed concept representation. They measure the change in the corresponding class’s logit score. CDISCO [14] has two measures. The first measure (M1) makes Constant Deletions in the pixel space and measures the number of classes that have a greater than 80% accuracy degradation. The second measure (M2) deletes concepts by zeroing out parameters in the model’s weight space that are associated with the concept, and they plot classification accuracy. MCD [35] deletes concepts in the pixel space by using inpainting. They measure model classification accuracy with respect to the fraction of inpainted pixels, unlike all prior measures that use the number of deleted concepts.
4.1.1 Limitations of Deletion-Based Measures
@@
There are two notable issues with deletion-based metrics that prevent them from accurately evaluating faithfulness. First, it is unclear how to delete a concept. Concepts are usually deleted by Constant Deletion, commonly with a baseline of 0. The baseline must entirely delete the presence of the concept without affecting other concepts. However, the feature attribution literature has shown that common baseline choices do not guarantee that a feature is completely deleted without affecting other features, leading to unfaithful explanations [11]. Removing features by marginalization is more accurate but requires evaluating expensive, high-dimensional expectations.
@@
Second, there are no guarantees that representations in the Perturbation Space will stay on manifold after deletion. The feature attribution literature [11, 2, 34, 33, 40] has found that the perturbed inputs used for creating explanations may be unrealistic, lying off the data manifold. In the off-manifold regions of the data space, a highly non-linear model behaves unpredictably, so the model’s outputs may not be meaningfully related to input features. Thus, measuring feature importance with off-manifold inputs can lead to incorrect explanations [11]. More seriously, these explanations can be adversarially attacked; in [33], a model’s off-manifold behavior was modified to hide its dependence on undesirable features used for on-manifold inputs. Such issues can be mitigated by generating explanations using on-manifold input perturbations but requires accurate modeling of the input distribution [11, 34, 36, 43, 15].
Deletions in the concept or weight space manipulate intermediate representations, so the same concerns hold. Thus, we argue that deletion metrics have significant unresolved issues that introduce unreliability in their faithfulness evaluations, so we avoid using these approaches.
@@ 4.2 Surrogate-Based Measures
Surrogate-based measures evaluate faithfulness by approximating (Eq 3). Any surrogate-based measure must make two choices: 1) what the surrogate $s$ is, and 2) what the metric $d$ is. We will introduce prior faithfulness measures, and then we discuss their limitations.
4.2.1 Desidarata
The faithfulness definition in Eq. 3 does not constrain the surrogate $s$, but there are certain properties our surrogate should have. First, $s$ should be as simple as possible. Recognize that $s$ reflects the mental computation any human interpreter may be expected to do when trying to map the explanation onto the model’s prediction. An interpretable explanation that requires a complex $s$ is effectively useless; it will not give the human interpreter any meaningful insight into how the model reaches its prediction. Second, the $s$ should incorporate all components of the explanation. For example, U-CBEMs explain with both $V_{i}$ and $A_{i}$, so both should be used by the surrogate. If a component is not included, its impact on faithfulness cannot be measured. Third, we desire metric(s) $d$ that rewards explanations that can closely reconstruct all model outputs (e.g. across all classes), instead of only the model’s predicted class. This directly follows from Eq. 3.
Table 2: Surrogate-Based Measures: We compare prior surrogate-based measures introduced in ICE and C-SHAP to the proposed Surrogate Faithfulness (SURF) measure. According to our desidarata, we require surrogates that are simple (e.g., not reconstruction-based and not requiring additional parameters), incorporate concept importances into the surrogate, and measure errors across all outputs. Prior works fail to meet these desidarata (except for ICE not requiring additional parameters). We mark in green and red the properties that meet our desidarata. A ✓  or ✗  denotes that the method does or does not have the property.
@@ 4.2.1 Desidarata
4.2.2 Prior Surrogate-Based Measures
Prior works have introduced two surrogate-based measures:
ICE-Eval: ICE introduces its own evaluation measure (termed ICE-Eval). ICE-Eval’s surrogate assumes the U-CBEM’s projection operation $P$ has a reconstruction function $\tilde{\mathcal{P}}^{-1}:\mathbb{R}^{K}\to\mathcal{H}$. Their surrogate then is:
$s_{i}(.)$ $=f_{i}(\tilde{\mathcal{P}}^{-1}(\mathcal{P}(\mathbf{h};V_{i})))\;\;\forall\;i$ (4)
where $f_{i}(.)$ denotes the $i$’th output from function $f(.)$. ICE-Eval introduces two metrics. Both are only applicable to classification models.
where $t$ is the index corresponding to the groundtruth class of $\mathbf{x}$. The second metric is:
@@ 4.2.2 Prior Surrogate-Based Measures
C-SHAP-Eval: C-SHAP introduces its own evaluation measure (termed C-SHAP-Eval). Its surrogate learns a two-layer perceptron $\psi_{\text{C-SHAP}}:\mathbb{R}^{K}\to\mathcal{H}$:
$s_{i}(.)$ $=f_{i}(\psi_{\text{C-SHAP}}(\mathcal{P}(\mathbf{h};V_{i})))\;\;\forall\;i$ (7)
The metric they define is:
where $a_{r}$ is the accuracy from random predictions.
ICE-Eval and C-SHAP-Eval fail all three of our desidarata. (1) A critical piece of both surrogates is to reconstruct the model’s original embedding, and this reconstruction may be non-linear. This step makes the explanation indirectly related to the model’s output, placing a larger burden on the human interpreter. Additionally, C-SHAP introduces learnable parameters via $\psi_{\text{C-SHAP}}$. (2) Neither surrogate depends on concept importances ($A_{i}^{\prime}s$). Thus, unfaithful $A_{i}$’s will not alter faithfulness scores. (3) Error is measured with respect to the predicted or groundtruth class, ignoring errors on the remainder of the output distribution.
@@ 5 Method: Surrogate Faithfulness (SURF)
Surrogate Faithfulness (SURF) has two components. A simple surrogate maps the explanation to the model’s output (Fig. 1, Step 3); then, outputs from the surrogate and the original model are compared (Fig. 1, Step 4).
@@ 5.1 The SURF Surrogate
The final linear layer $\mathbf{F}=\begin{bmatrix}\mathbf{F}_{1}&\cdots&\mathbf{F}_{C}\end{bmatrix}$ operates on representation $\mathbf{H}$ to give the output $\mathbf{y}\in\mathcal{Y}$. In terms of their components, $\mathbf{H}=\begin{bmatrix}\mathbf{h}_{1}&\cdots&\mathbf{h}_{HW}\end{bmatrix}^{T}$ and $\mathbf{F}_{i}=\begin{bmatrix}\mathbf{f}_{i,1}&\cdots&\mathbf{f}_{i,HW}\end{bmatrix}^{T}\in\mathbb{R}^{(HW)\times D}$. Thus, output $y_{i}$ of $\phi$ is given by:
@@
We define SURF’s surrogate $s$ to take the form of the linear layer (Eq. 9), where we input the concept basis representation and $A_{i}$’s from any U-CBEM. The surrogate is:
$\hat{y_{i}}$ $=\sum_{j=1}^{HW}\sum_{k=1}^{K}\alpha_{i,k}\mathcal{P}(\mathbf{h}_{j};V_{i})_{k}\quad\forall\ i$ (10)
where $\mathcal{P}(.)_{k}$ denotes the $k$’th element. Following prior U-CBEMs and evaluation measures, SURF also operates only on the model’s final linear layer, allowing its linear surrogate to fully represent the model’s computation.
Table 3: Measure-over-measure comparison. We compare surrogate-based measures across three explanation settings (Perfect, Rand Imp, and Full Rand). We expect evaluations in the Perfect setting to give perfect faithfulness scores, and we expect progressively worse faithfulness evaluations as we increase randomness (just importances in Rand Imp and fully random in Full Rand). We include the SURF metrics (MAE and EMD) along with other metrics used by prior works. The SURF measure behaves as expected, whereas C-SHAP-Eval reports higher faithfulness in the Full Rand setting and ICE-Eval reports perfect faithfulness scores in the Rand Imp setting.
Surrogate Top-1 (%) Rank Corr Norm L1 MAE EMD Learnable Params.
Perfect C-SHAP-Eval (CEL) 9.06 -0.18 0.94 1.72 0.842 1M
C-SHAP-Eval (L1) 0.51 0.44 1.34 0.86 0.846 1M
ICE-Eval 100 1.00 0.00 0.00 0.000 0
[…]
Rand Imp C-SHAP-Eval (CEL) 9.06 -0.18 0.94 1.72 0.842 1M
C-SHAP-Eval (L1) 0.51 0.44 1.34 0.86 0.846 1M
ICE-Eval 100 1.00 0.00 0.00 0.000 0
[…]
Full Rand C-SHAP-Eval (CEL) 86.9 0.42 12.2 14.9 0.249 1M
C-SHAP-Eval (L1) 7.7 0.53 1.29 0.729 0.853 1M
ICE-Eval 3.4 0.07 1.00 1.66 0.862 0
[…]
Figure 1: Surrogate Faithfulness (SURF) overview. (1) Probe the model with a dataset of inputs to obtain intermediate embeddings from the penultimate layer $L$. (2) A training subset of the embeddings is used to fit the U-CBEM, discovering CAVs and concept importances for all classes $C$. (3) A test set of embeddings (along with the learned U-CBEM parameters) are passed to the surrogate to be mapped to predictions. (4) The SURF metrics compare surrogate and model outputs to measure faithfulness.
@@ 5.2 The SURF Metrics
The SURF metrics measure the difference in model and surrogate outputs (Fig. 1, Step 4). For classification models, this means the logits. However, a model’s logit-space is unconstrained, varying drastically between models. This makes logit-space metrics difficult to interpret. Furthermore, metrics in the logit-space dilute the importance of the predicted class by aggregating errors over all classes. On the other hand, the probability-space (after softmax) normalizes the logits, allowing for a bounded, interpretable metric. However, low error in the probability-space does not always imply low error in the logit-space. Due to the softmax, the probability-space emphasizes the predicted class and diminishes the other classes, so one could achieve a low error by accurately reproducing the predicted class and ignoring the others. Since each space addresses the other’s flaws, SURF measures errors in both spaces.
@@
Let $\phi$ be a classification model, where $\mathbf{y}\in\mathbb{R}^{C}$ denotes class logits. To measure errors in the logit-space, we use the mean absolute error (MAE) between the logits:
$MAE=\frac{1}{|\mathcal{V}|C}\sum_{\mathbf{x}\in\mathcal{V}}\sum_{i=1}^{C}|y_{i}-\hat{y}_{i}|$ (11)
@@
where $\mathcal{V}$ is the set of inputs used to evaluate for faithfulness. To measure errors in the probability-space, we find the Earth Mover’s Distance (EMD) using a constant distance cost. Letting $\mathbf{p}\in\mathbb{R}^{C}$ denote class probability scores,
$EMD=\frac{1}{2|\mathcal{V}|}\sum_{\mathbf{x}\in\mathcal{V}}\sum_{i=1}^{C}|p_{i}-\hat{p}_{i}|$ (12)
@@
For a U-CBEM to be faithful, its explanations must have low EMD and MAE errors. For regression, we only use MAE.
@@
In contrast to prior surrogate-based approaches, Surrogate Faithfulness (SURF) meets all of our desidarata. (1) It introduces a simple surrogate with no additional learnable parameters: the output for any class $i$ is a linear combination of the concept basis, weighted by the concept importances ($A_{i}$). The surrogate is not reconstruction-based; it tries to predict the next linear representation (e.g., class logits), instead of reconstructing embedding $\mathbf{h}$. Thus, only concepts useful for the prediction are needed, which may only be a subset of the concepts needed for reconstructing $\mathbf{h}$ from $\mathcal{P}(\mathbf{h};V)$. (2) SURF’s surrogate incorporates $A_{i}$, allowing inaccuracies in $A_{i}$ to affect faithfulness scores. (3) SURF defines two metrics $d$ that measure errors across the entire output space, instead of specific classes, to comprehensively assess faithfulness. These differences are summarized in Tab. 2.
6 Experiments
We perform experiments to: (1) empirically prove that SURF is a better faithfulness measure than prior ones; (2) benchmark prior U-CBEMs using SURF across three tasks, finding that many state-of-the-art U-CBEMs have poor faithfulness; (3) analyze the impact of the number of concepts on faithfulness; (4) study the robustness of U-CBEMs in adversarial settings, a novel scenario that has been unexplored in prior evaluations.
@@ 6 Experiments
Table 4: Benchmarking U-CBEM faithfulness. We apply SURF to evaluate explanations from prior U-CBEMs in three tasks. Along with the SURF metrics (MAE and EMD), we report other metrics to serve as a comparison. We find that most prior U-CBEMs are not faithful, as indicated by large errors in the logit and probability space.
@@
U-CBEM MAE ($\downarrow$) EMD ($\downarrow$) Top-1 (%) ($\uparrow$) Rank Corr ($\uparrow$)
CDISCO 2.29 0.957 0.9 0.000
ICE 3.06 0.438 93.1 0.222
CRAFT 1.70 0.844 78.9 0.484
C-SHAP 2.40 0.587 41.7 0.461
MCD 1.42 0.079 97.0 0.808
@@
MAE ($\downarrow$) Attr-Acc (%) ($\uparrow$)

### v3

@@ 4.1 Deletion-Based Proxies
Deletion Method: How concepts are removed from the Deletion Space, usually by setting the concept to a baseline value (e.g., 0).
@@
Except for CRAFT, proxies set the concept projection to identity and the surrogate to $f(.)$ from the original model. Prior proxies (organized in Tab. 1) make different choices along each italicized factor mentioned above. ACE deletes concepts in the image space by setting associated pixels to a constant value (Constant Deletion). They measure the model’s classification accuracy after each deletion across a dataset. MCD measures classification accuracy after deleting concepts in the pixel space with inpainting. HU-MCD similarly measures classification accuracy but performs deletions through a masking strategy which simulates running the model on the irregularly-shaped input, better ignoring masked regions. CDISCO has two measures. The first measure (M1) makes Constant Deletions on pixels and measures the number of classes whose accuracy degrades by more than 80%. The second measure (M2) deletes concepts by zeroing out model parameters associated with the concept, and they report classification accuracy. CRAFT uses the U-CBEMs defined $\mathcal{P}$ to transform $\mathbf{h}$ to a concept representation. It makes Constant Deletions in the concept space by setting the concept to zero; then, it reconstructs $\mathbf{h}$ from the perturbed concept representation and measures the change in the corresponding class’s logit score.
@@
There are two notable issues with deletion-based proxies that prevent them from accurately evaluating faithfulness. First, it is unclear how to delete a concept. Concepts are usually deleted by Constant Deletion, commonly with a baseline of 0. This baseline must entirely delete the presence of the concept without affecting other concepts. However, the feature attribution literature has shown that common baseline choices do not guarantee that a feature is completely deleted without affecting other features, leading to unfaithful explanations [Frye2020ShapleyManifold]. Removing features by marginalization is more accurate but requires evaluating expensive, high-dimensional expectations.
@@
Second, there are no guarantees that representations in the Deletion Space will stay on manifold after deletion. The feature attribution literature [Frye2020ShapleyManifold, akata2022manifold, taufiq2023manifold, Slack2019FoolingLA, yeh2022threading] has found that the perturbed inputs used for creating explanations may be unrealistic, lying off the data manifold. In the off-manifold regions of the data space, a highly non-linear model behaves unpredictably, so the model’s outputs may not be meaningfully related to input features. Thus, calculating feature importance with off-manifold inputs can lead to incorrect explanations [Frye2020ShapleyManifold]. More seriously, these explanations can be adversarially attacked; in [Slack2019FoolingLA], a model’s off-manifold behavior was modified to hide its dependence on undesirable features used for on-manifold inputs. Such issues can be mitigated by using on-manifold input perturbations, but this requires accurate modeling of the input distribution [Frye2020ShapleyManifold, taufiq2023manifold, vu2022emap, zaher2024manifold, Hase2021TheOP]. These problems are not unique to feature attribution; they extend to deletions in the concept and weight space, which make similar manipulations in the model’s intermediate representation space.
Thus, we argue that deletion-based proxies are unreliable, and we avoid using or benchmarking against them.
@@ 4.2 Surrogate-Based Measures
Any surrogate-based measure must choose the surrogate $s$ and the metric $d$. For these measures, the input image, and thus the intermediate representation, is the same in both paths of Fig. 1 (left), and the concept projection is defined by the U-CBEM being evaluated.
4.2.1 Desiderata
The faithfulness definition in Eq. 3 does not constrain the surrogate $s$, but there are certain desiderata our surrogate should satisfy. First, $s$ should be as simple as possible. Recognize that $s$ reflects the mental computation any human interpreter is expected to do when trying to connect the explanation to the model’s prediction. An interpretable explanation that requires a complex $s$ just shifts the interpretation issue downstream; the human interpreter still lacks insight into how the model reaches its prediction. Second, $s$ should incorporate all components of the explanation. For example, U-CBEMs explain with both $V_{i}$ and $A_{i}$, so both should be used by $s$. If a component is not included, its impact on faithfulness cannot be measured. Third, we desire metric(s) $d$ that rewards explanations that closely reconstruct all of $\mathbf{y}$ (e.g. across all classes), not just specific (e.g., predicted) classes.
4.2.2 Prior Surrogate-Based Measures
Two prior works have introduced surrogate-based measures:
ICE-Eval: ICE introduces its own evaluation measure (we term as ICE-Eval). ICE-Eval’s surrogate assumes the U-CBEM’s projection operation $P$ has a reconstruction function $\tilde{\mathcal{P}}^{-1}:\mathbb{R}^{K}\to\mathcal{H}$. Their surrogate then is:
$\mathbf{\hat{y}}_{i}=s_{i}(.)$ $=f_{i}(\tilde{\mathcal{P}}^{-1}(\mathcal{P}(\mathbf{h};V_{i})))\;\;\forall\;i$ (4)
where $f_{i}(.)$ denotes the $i$’th output from function $f(.)$. ICE-Eval uses two metrics (only applicable to classification):
where $t$ is the index corresponding to $\mathbf{x}$’s groundtruth class and $\mathcal{V}$ is the set of inputs used to evaluate faithfulness. Letting $\mathbf{p}$ denote class probability scores, the second metric is:
[…]
C-SHAP-Eval: C-SHAP introduces its own evaluation measure (we term as C-SHAP-Eval). Its surrogate learns a two-layer perceptron $\psi_{\text{C-SHAP}}:\mathbb{R}^{K}\to\mathcal{H}$:
$\mathbf{\hat{y}}_{i}=s_{i}(.)$ $=f_{i}(\psi_{\text{C-SHAP}}(\mathcal{P}(\mathbf{h};V_{i})))\;\;\forall\;i$ (7)
The metric they define is:
where $a_{r}$ is the accuracy from random predictions.
ICE-Eval and C-SHAP-Eval fail all three of our desiderata. (1) A critical piece of both surrogates is to reconstruct the model’s original embedding, and this reconstruction may be non-linear. Additionally, C-SHAP introduces non-linearity and learnable parameters via $\psi_{\text{C-SHAP}}$. The overly complex surrogates place a large burden on the human interpreter, which is not reflected in prior interpretability and faithfulness evaluations. (2) Neither surrogate depends on concept importances ($A_{i}^{\prime}s$). Thus, unfaithful $A_{i}$’s will not alter faithfulness scores. (3) Error is measured only on the predicted or groundtruth class, ignoring errors on the remainder of the output distribution.
Table 2: Surrogate-based desiderata We compare prior surrogate-based measures introduced in ICE and C-SHAP to the proposed Surrogate Faithfulness (SURF) measure. According to our desiderata, we require surrogates that are simple (e.g., not reconstruction-based and not requiring additional parameters), incorporate concept importances into the surrogate, and measure errors across all outputs. Prior measures largely fail to meet these desiderata. We mark in green and red the properties that meet our desiderata. A ✓  or ✗  denotes that the method does or does not have the property.
@@ 4.2.2 Prior Surrogate-Based Measures
Table 3: Measure-over-measure comparison. We compare surrogate-based measures across three explanation settings (Perfect, Rand Imp, and Full Rand). We expect evaluations in the Perfect setting to give perfect faithfulness scores, and we expect progressively worse faithfulness evaluations as we increase randomness (just importances in Rand Imp and fully random in Full Rand). We include $\text{SURF}_{\text{EMD}}$ and $\text{SURF}_{\text{MAE}}$ along with metrics used by prior works. The SURF metrics behaves as expected, whereas C-SHAP-Eval reports higher faithfulness in the Full Rand setting and ICE-Eval reports perfect faithfulness scores in the Rand Imp setting.
Surrogate Top-1 (%) ($\uparrow$) Rank Corr ($\uparrow$) Norm L1 ($\downarrow$) $\text{{SURF}}_{\text{MAE}}$ ($\downarrow$) $\text{{SURF}}_{\text{EMD}}$ ($\downarrow$) Params Learnt ($\downarrow$) FLOPs ($\downarrow$)
Perfect C-SHAP-Eval (CEL) 9.02 -0.02 1.27 1.97 0.865 1M 205M
C-SHAP-Eval (L1) 6.13 0.08 2.15 0.54 0.883 1M 205M
ICE-Eval 100 1.00 0.00 0.00 0.000 0 614K
@@
Rand Imp C-SHAP-Eval (CEL) 9.02 -0.02 1.27 1.97 0.865 1M 205M
C-SHAP-Eval (L1) 6.13 0.08 2.15 0.54 0.883 1M 205M
ICE-Eval 100 1.00 0.00 0.00 0.000 0 614K
[…]
Full Rand C-SHAP-Eval (CEL) 97.6 0.02 181.7 168.2 0.125 1M 205M
C-SHAP-Eval (L1) 6.1 0.08 3.59 1.721 0.883 1M 205M
ICE-Eval 3.3 0.00 1.00 3.17 0.882 0 614K
[…]
@@ 5 Method: Surrogate Faithfulness (SURF)
Surrogate Faithfulness (SURF) introduces a simple, linear surrogate and two metrics; taken together, they allow SURF to satisfy the desiderata.
@@ 5.1 The SURF Surrogate
The final linear layer $\mathbf{F}=\begin{bmatrix}\mathbf{F}_{1}&\cdots&\mathbf{F}_{C}\end{bmatrix}$ operates on representation $\mathbf{H}$ to give the output $\mathbf{y}\in\mathcal{Y}$. In terms of their components, $\mathbf{H}=\begin{bmatrix}\mathbf{h}_{1}&\cdots&\mathbf{h}_{HW}\end{bmatrix}^{T}$ and $\mathbf{F}_{i}=\begin{bmatrix}\mathbf{f}_{i,1}&\cdots&\mathbf{f}_{i,HW}\end{bmatrix}^{T}\in\mathbb{R}^{(HW)\times D}$. Thus, output $y_{i}$ of $\phi$ is given by (up to the bias term):
@@
We define SURF’s surrogate $s$ to take the form of the linear layer (Eq. 9), where we use the concept representation and $A_{i}$’s obtained from any U-CBEM. The surrogate is:
$\mathbf{\hat{y}}_{i}$ $=s_{i}(.)=\sum_{j=1}^{HW}\sum_{k=1}^{K}\alpha_{i,k}\mathcal{P}(\mathbf{h}_{j};V_{i})_{k}\quad\forall\ i$ (10)
where $\mathcal{P}(.)_{k}$ denotes the $k$’th element. Following prior U-CBEMs and evaluation measures, SURF also operates only on the model’s final linear layer. Thus, its surrogate fully represents the model’s computation, while greatly reducing complexity compared to prior faithfulness surrogates.
@@ 5.2 The SURF Metrics
The SURF metrics measure the difference in model and surrogate outputs. For classification models, this means the logits. However, a model’s logit-space is unconstrained, varying drastically between models. This makes logit-space metrics difficult to interpret. Furthermore, metrics in the logit-space dilute the importance of the predicted class by aggregating errors over all classes. On the other hand, the probability-space (after softmax) normalizes the logits, allowing for a bounded, interpretable metric. However, low error in the probability-space does not always imply low error in the logit-space. Due to the softmax, the probability-space emphasizes the predicted class and diminishes the other classes, so one could achieve a low error by accurately reproducing the predicted class and ignoring the others. Since each space addresses the other’s flaws, SURF measures errors in both spaces.
@@
If $\phi$ is a classification model, then $\mathbf{y}$ denotes class logits. To measure errors in the logit-space, we use the mean absolute error between the logits:
$\text{SURF}_{\text{MAE}}=\frac{1}{|\mathcal{V}|C}\sum_{\mathbf{x}\in\mathcal{V}}\sum_{i=1}^{C}|y_{i}-\hat{y}_{i}|$ (11)
@@
To measure errors in the probability-space, we find the Earth Mover’s Distance using a constant distance cost.
$\text{SURF}_{\text{EMD}}=\frac{1}{2|\mathcal{V}|}\sum_{\mathbf{x}\in\mathcal{V}}\sum_{i=1}^{C}|p_{i}-\hat{p}_{i}|$ (12)
@@
A faithful U-CBEM will have explanations with low $\text{SURF}_{\text{EMD}}$ and $\text{SURF}_{\text{MAE}}$ errors. For regression, we only use $\text{SURF}_{\text{MAE}}$.
@@
In contrast to prior surrogate-based measures, SURF meets all of our desiderata. (1) It introduces a simple surrogate with no additional learnable parameters: the output for any class $i$ is a linear combination of the concept representation, weighted by the concept importances ($A_{i}$). The surrogate is not reconstruction-based; it tries to linearly predict the next representation (e.g., class logits), instead of reconstructing embedding $\mathbf{h}$. Thus, only concepts useful for the prediction are needed, which may be a subset of the concepts needed for reconstructing $\mathbf{h}$ from $\mathcal{P}(\mathbf{h};V)$. (2) SURF’s surrogate incorporates $A_{i}$, so inaccuracies in $A_{i}$ impact faithfulness scores. (3) SURF defines two metrics $d$ that measure errors across the entire output space, instead of specific classes, to comprehensively assess faithfulness. These differences are summarized in Tab. 2.
@@
Table 4: Benchmarking U-CBEM faithfulness. We apply SURF to evaluate explanations from prior U-CBEMs in three tasks. Along with the $\text{SURF}_{\text{MAE}}$ and $\text{SURF}_{\text{EMD}}$, we report other metrics to serve as a comparison. We find that prior U-CBEMs are not faithful, as indicated by large errors in the logit and probability space.
@@
U-CBEM $\text{{SURF}}_{\text{MAE}}$ ($\downarrow$) $\text{{SURF}}_{\text{EMD}}$ ($\downarrow$) Top-1 (%) ($\uparrow$) Rank Corr ($\uparrow$)
CDISCO 3.40 0.932 0.2 0.002
ICE 3.33 0.628 98.9 0.093
CRAFT 3.19 0.878 90.6 0.068
C-SHAP 3.28 0.882 6.3 0.005
MCD 2.60 0.426 99.4 0.145
HU-MCD 1.97 0.384 99.7 0.149
@@
$\text{{SURF}}_{\text{MAE}}$ Attr-Acc (%) ($\uparrow$)
@@
6 Experiments
We first perform a measure-over-measure study with a basic sanity check: do faithfulness scores decrease as we progressively randomize the explanation? Only SURF passes this check. Then, we benchmark prior U-CBEMs across three tasks, revealing that SOTA methods are not faithful. Finally, we leverage SURF to analyze the trade-off between the number of concepts in an explanation and faithfulness.